5^2 + 12^2 = 13^2
25 + 144 = 169
169 = 169 I understand the mathwork involved in Pathagorean Triples and Theorems, but I am having trouble with my assignment. The assignment says to find a method or formula outside of my textbook and generate five pathagorean triples. I need to confirm validity of these pythagorean triples using the pythagorean theorem formula: a2 + b2 = c2, but Please remember that a2 + b2 = c2 is NOT the formula you're looking for to generate the triples with, it is only to verify each of the
triples (at least 5). I am confused plz help me.

Answer by scott8148(6628)

a^2 + b^2 - 2b^2 = a^2 - b^2 = (a + b)(a - b)

a^2 + b^2 - 2b^2 = c^2 - 2b^2

(a + b)(a - b) = c^2 - 2b^2 ___ (a + b)(a - b) + 2b^2 = c^2

the sum times the difference, plus twice the smaller squared; equals a perfect square


this will work, but I'm sure there must be other methods